OpenStudy (anonymous):

I need help on : A.) 1 + a/5 = -1 (A=?) B.) y/2 + 5 = -12 (Y=?) C.) x/3 - 9 = 0 (X=?) D.) 14 + h/5 = 2 (H=?) E.) m/8 + 4 = 16 (M=?) F.) a/4 - 21 = 7 (A=?) G.) 41 = 2/5x - 7 (X=?) H.) -3 + m/3 = 12 (M=?) Please show work for me to understand!

OpenStudy (alexwee123):

A) 1 + a/5 = -1 so the point of this problem is to isolate "a" so you would subtract 1 from both sides... 1+a/5 -1=-1-1 a/5=-2 then just multiply both sides by 5 to get rid of the denominator under "a" 5(a/5)=-2(5) a=-10 can you try some of the other problems by yourself?